OpenStudy (anonymous):

R(t) = 10 sin (30t) + 10 How would I graph this?

OpenStudy (turingtest):

\[y=A\sin(Bx+C)\]has amplitude A, period \[\frac{2\pi}{|B|}\]and phase shift C the graph of \[f(x)+k\]is the graph of f(x) shifted up by k units so your graph is a sin wave with an amplitude of 10, a period of pi/15, and a phase shift of zero, moved upwards by 10 units http://www.wolframalpha.com/input/?i=plot+y+%3D+10+sin+%2830t%29+%2B+10

OpenStudy (anonymous):

for period, can't I just do 360/k?

OpenStudy (turingtest):

\[2\pi\text{radians}=360^{\circ}\]so they are the same. radians is better though, so get used to them

OpenStudy (turingtest):

and if you mean what I called k, then no it would be 360/B for the period

OpenStudy (anonymous):

so it'll be 30?

OpenStudy (turingtest):

what will be 30?

OpenStudy (anonymous):

wait, how would I find the period? would it be 360/30? and would the maximum value be positive 10?